in a chem class, 7 liters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. How many liters are needed?
I tried : to make a chart
or .10x + .04{7 + X}= .06
I got a neg. number. Please help.

Answer by Paul(988)

Let the solution be x:
7(4)-10x=6(x-4)
28-10x=6x-24
16x=52
x=3.25
Hence, about 3.25L is needed.
PAul.